OpenStudy (anonymous):

endpoints of a diameter (-7,8),(-3,-2). Find the equation for the circle

OpenStudy (anonymous):

I need some help to figure this out.

OpenStudy (anonymous):

Well you need origin and the radius so you can plug it in to the equation of a circle To find the origin use the midpoint theorem... To find the radius use distance formula of the points and divide by two. Need me to evaluate?

OpenStudy (anonymous):

Yes please.

OpenStudy (anonymous):

Formula of a circle in standard form... (x-h)^2+(y-k)^2=r^2... Where (h,K) is the origin and r is the radius.. Midpoint formula is \[(\frac{ x _{1} +x_2 }{ 2 }, \frac{ y_1+y_2 }{ 2 })\] Distance Formula is http://cs.selu.edu/~rbyrd/math/distance/ (too lazy to type it out

OpenStudy (anonymous):

So I came up with 8. Is that correct?

OpenStudy (anonymous):

the anser is the whole equation?

OpenStudy (anonymous):

Ok. I got 5 for the x and 3 for y. So the equation should be (x+5)+(y-3).

OpenStudy (anonymous):

=29